# BIBO stability of port-Hamiltonian systems

### v1

@@ Abstract.
We study the question of when a distributed port-Hamiltonian system is bounded-input bounded-output (BIBO) stable. Exploiting the particular structure of the transfer function of these systems, we derive several sufficient conditions for BIBO stability.
@@ 1. Introduction
Since their introduction port-Hamiltonian systems (pHS) have constituted an important system class used in modelling and control of real world applications. Their close connection to the ideas of energy-flow and energy-conservation makes them a useful tool in the study of physical systems [27, 28, 6].
A particular type of pHS are so-called distributed port-Hamiltonian systems [29, 13, 12, 30]. These contain systems formally given by a PDE of the form
@@
on some interval $[a,b]$ together with boundary control and observation at its endpoints. The justification and motivation for the study of theses systems lies in their ability to provide a modelling framework for a variety of different real-world processes, including first of all transport and flow phenomena such as in fluid mechanics [17] or chemistry [34, 33] applications and also vibrations within flexible structures as they appear e.g. in robotics [15, 14]. In particular in this latter case, the approach using infinite-dimensional systems allows to produce refinements of finite-dimensional models where the mentioned phenomena (vibrations, flows, etc. ) are usually ignored or only treated in a coarse way. We refer the interested reader to the literature review [18] for a more extensive look on the history, study and application of distributed pHS.
@@
Only recently work has been undertaken to systematically classify different notions of stability for distributed port-Hamiltonian systems. These studies considered in particular exponential [24] and asymptotic stability [32]. Previously only some sufficient conditions for these properties where known [12, 30], which were however mainly restricted to the case of systems featuring a contraction semigroup as encountered in the vast majority of concrete examples.
@@
Besides these different internal stability concepts, the input-output behaviour of pHS is an important aspect. It is a nice feature that a distributed pHS – under practically weak technical assumptions – is $L^{2}$-well-posed [12]. Thus in particular there exist constants $c_{t}$ for any $t>0$ such that for any input $u\in L^{2}\left([0,t],U\right)$ the corresponding output $y$ satisfies $\|y\|_{L^{2}\left([0,t],Y\right)}\leq c_{t}\|u\|_{L^{2}\left([0,t],U\right)}$. The importance of this result is again grounded in the port-Hamiltonian structure being used to model energy-flows and the input-output stability thus being related to the concepts of energy-conserving and energy-dissipating systems.
@@
In this contribution we study the notion of bounded-input bounded-output stability (BIBO stability), which is classical for finite-dimensional systems, for distributed port-Hamiltonian systems. An input-output system is called BIBO stable if any bounded input is mapped to a bounded output and there exists a uniform relative norm bound [26, 4, 31, 5, 1, 21, 9].
The question of whether a system is BIBO stable or not, while relatively straightforward in the finite-dimensional setting, becomes significantly more involved in the case of an infinite-dimensional state space such as for the pHS considered here. For example, it is then no longer the case that BIBO stability is always implied by exponential stability (see [21, Thm. 5.1] but note that crucially this example is not of port-Hamiltonian form).
In [21] it was shown that, for a linear system described in the framework of system nodes (see [23]) with finite-dimensional input and output spaces — as will be the case in the system class considered in this contribution — BIBO stability is equivalent to the inverse Laplace transform of the transfer function $\mathbf{G}(s)$ being a measure of bounded total variation. This includes in particular the case that the impulse response – upon being well-defined – of the system is an $L^{1}$ function.
@@
While this condition offers a complete characterisation of when a system is BIBO stable, it is often of limited applicability for concrete systems. Typically no closed form expression for the inverse Laplace transform of the respective transfer function exists or is prohibitively hard to find. Thus BIBO stability is in general much more difficult to establish than e.g. $L^{2}$-well-posedness. This circumstance justifies the search for other sufficient or necessary conditions and the concrete study of BIBO stability for particular system classes. A few such results have been provided in for example [21] already, however almost exclusively for parabolic systems, which exclude pHSs as considered here.
@@
We note that BIBO stability — besides being an intriguing topic in itself and classical control theory — plays an essential role in the applicability of funnel control to relative degree systems [10, 2]. Such systems with interior dynamics modelled by port-Hamiltonian systems appear for example in the study of mechanical systems containing flexible and deformable parts [15], which significantly influence the behaviour.
@@
This paper begins with a recap of the definition of a class of distributed port-Hamiltonian systems in Section 2. Then, in Section 3, we derive an abstract decomposition of the transfer function and exploit its particular structure in order to apply the known characterisation of BIBO stability in a general setting. This yields a collection of sufficient criteria expressed in the form of matrix conditions for a subclass of pHS. Section 4 then presents a different approach to BIBO without any reference to transfer functions, instead employing the $L^{1}$-well-posedness of pHS on an $L^{1}$ state space and the connection between $L^{1}$ and $L^{\infty}$ stability notions discussed in [20].
@@ 1.1. Notation
Let $A$ be the generator of a strongly continuous semigroup $\mathbb{T}=\left(\mathbb{T}(t)\right)_{t\geq 0}$ on a Banach space $X$. $X_{1}$ is defined as the space $\mathrm{D}(A)$ with the norm $\|x\|_{X_{1}}:=\|(\beta I-A)x\|_{X}$ with $\beta\in\rho(A)$ and $X_{-1}$ as the completion of $X$ with respect to the norm $\|x\|_{X_{-1}}:=\|(\beta I-A)^{-1}x\|_{X}$ again with $\beta\in\rho(A)$. There exists a unique extension $\mathbb{T}_{-1}=\left(\mathbb{T}_{-1}(t)\right)_{t\geq 0}$ of the semigroup $\mathbb{T}$ to the space $X_{-1}$ with generator $A_{-1}:X\rightarrow X_{-1}$ which is an extension of the operator $A$. For more details see [23, 25].
For a matrix $M\in\mathbb{K}^{n\times n}$ and $p,q\in[1,\infty]$ let $\|M\|_{\ell^{p}\rightarrow\ell^{q}}$ be the matrix norm defined with respect to the $\ell^{p}$-norms on $\mathbb{K}^{n}$ given by $\|x\|_{\ell^{p}}=\left(\sum_{k=1}^{n}(x_{k})^{p}\right)^{\frac{1}{p}}$ for $p\in[1,\infty)$ and $\|x\|_{\ell^{\infty}}=\sup_{1\leq k\leq n}|x_{k}|$ for $p=\infty$ as
@@
For a Borel measure $h$ on $[0,\infty)$, let $\mathcal{L}\left\{h\right\}$ denote its Laplace transform if it exists (on some right half-plane). For a function $g$ of one complex variable, if there exists a measure $\mu$ such that $\mathcal{L}\left\{\mu\right\}$ and $g$ agree on some right half plane we will call $\mu$ the inverse Laplace transform and write $\mu=\mathcal{L}^{-1}\left\{g\right\}$.
Let $\mathcal{M}([0,\infty),\mathbb{K}^{n\times m})$ denote the set of Borel measures of bounded total variation on $[0,\infty)$ with values in $\mathbb{K}^{n\times m}$. Furthermore, for $h\in\mathcal{M}([0,\infty),\mathbb{K}^{n\times m})$ let $\|h\|_{\mathcal{M}}$ denote the total variation of $h$ [8, Sec. 3.2].
2. Distributed port-Hamiltonian systems
In this work we consider linear distributed port-Hamiltonian systems [12] of the form
(2.1) $\frac{\partial x}{\partial t}(\xi,t)$ $=P_{1}\frac{\partial}{\partial\xi}\left(\mathcal{H}(\xi)x(\xi,t)\right)+P_{0}(\xi)\left(\mathcal{H}(\xi)x(\xi,t)\right),$ $(\xi,t)$ $\in[a,b]\times[0,\infty),$
@@ 2. Distributed port-Hamiltonian systems
with the following assumptions
@@
  • $\mathcal{H}\in L^{\infty}\left([a,b],\mathbb{K}^{n\times n}\right)$, $\mathcal{H}(\xi)$ is self-adjoint for almost all $\xi\in[a,b]$ and $\exists M,m>0$ s.t. $mI\leq\mathcal{H}(\xi)\leq MI$ for almost all $\xi\in[a,b]$;
  • $\widetilde{W}_{B}\in\mathbb{K}^{n\times 2n}$ and $\left[\begin{smallmatrix}\widetilde{W}_{B}\\ \widetilde{W}_{C}\end{smallmatrix}\right]\in\mathbb{K}^{2n\times 2n}$ have full rank.
We will in the following use $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ to denote this pHS.
In light of most of the pHS literature where contractive systems are considered, we explicitly state that we are making only the following weaker assumption.
Assumption 2.1.
@@
is the infinitesimal generator of a $C_{0}$-semigroup on the state space $X=L^{2}\left([a,b],\mathbb{K}^{n}\right)$ with inner product $\left\langle f,g\right\rangle_{X}:=\frac{1}{2}\int_{a}^{b}g(\xi)^{*}\mathcal{H}(\xi)f(\xi)\,\mathrm{d}\xi$.
Remark 2.2.
@@
    As the term $P_{0}\mathcal{H}$ represents merely a bounded perturbation within the operator $A$ that does not change its semigroup generation properties, Assumption 2.1 is in fact independent of the choice of $P_{0}$.
@@
    Contrary to the setting usually considered in the literature, we allow $P_{0}$ to depend on $\xi$. Note that the majority of results found e.g. in [12] for pHS with constant $P_{0}$ hold analogously in the case considered here.
@@
    The assumption of $A$ generating a $C_{0}$-semigroup is in particular satisfied if
    (2.2) $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}\geq 0$
    where $R_{0}=\frac{1}{\sqrt{2}}\left[\begin{smallmatrix}P_{1}&-P_{1}\\ \mathbbm{1}_{n}&\mathbbm{1}_{n}\end{smallmatrix}\right]$ and $\Sigma=\left[\begin{smallmatrix}0&\mathbbm{1}_{n}\\ \mathbbm{1}_{n}&0\end{smallmatrix}\right]$. If in addition $P_{0}^{\ast}=-P_{0}$, then (2.2) is even equivalent to the semigroup being contractive [12, Thm. 7.2.4, Thm. 10.3.1].
For a pHS satisfying Assumption 2.1,direct extension of [12, Thm. 11.3.2 & Thm. 12.1.3] to the case of spatially dependent $P_{0}$, shows that its transfer function on $\rho(A)$ can be found as the unique solution of
(2.3) $\begin{split}&sx_{0}(\xi)=P_{1}\frac{\mathrm{d}}{\mathrm{d}\xi}\left(\mathcal{H}(\xi)x_{0}(\xi)\right)+P_{0}(\xi)\left(\mathcal{H}(\xi)x_{0}(\xi)\right),\\ &u_{0}=\widetilde{W}_{B}\begin{bmatrix}\left(\mathcal{H}x\right)(b)\\ \left(\mathcal{H}x\right)(a)\end{bmatrix},\qquad\mathbf{G}(s)u_{0}=\widetilde{W}_{C}\begin{bmatrix}\left(\mathcal{H}x\right)(b)\\ \left(\mathcal{H}x\right)(a)\end{bmatrix}.\end{split}$
2.1. BIBO stability of distributed pHS
In this paper we want to study the question of BIBO stability of such distributed port-Hamiltonian systems.
In the following we will understand the term classical solution in the common sense of evolution equations, that is as a collection $(u,x,y)$ with $u\in C([0,\infty),U)$, $x\in C^{1}([0,\infty),X)$ and $y\in C([0,\infty),Y)$ solving Equation (2.1).
Definition 2.3.
A pHS is called BIBO stable if there exists $c>0$ such that for any of its classical solution $(u,x,y)$ with $u\in C^{\infty}_{c}([0,\infty),U)$ and $x(0)=0$ and for all $t>0$ we have that $\|y\|_{L^{\infty}\left([0,t],Y\right)}\leq c\|u\|_{L^{\infty}\left([0,t],U\right)}$.
Now, as any pHS satisfying Assumption 2.1 can be equivalently expressed within the framework of system nodes (see [23] for a general introduction and [12, Ch. 11][19] for the rewriting of pHS respectively general boundary control systems in this form), they fall within the class of systems studied in [21]. Hence, by Theorem 3.4 therein, BIBO stability is equivalent to the transfer function $\mathbf{G}$ being the Laplace transform of a measure of bounded total variation.
Remark 2.4.
@@ 2.1. BIBO stability of distributed pHS
    As any pHS can be rewritten as a system node and following the analogous construction in this formalism [23], we can define a generalised solution $(u,x,y)$ of a pHS for any input function $u\in L^{1}_{\mathrm{loc}}\left([0,\infty),U\right)$ and initial value $x_{0}\in X$.
    Here the state is given by $x(t)=\mathbb{T}(t)x_{0}+\int_{0}^{t}\mathbb{T}_{-1}\left(t-s\right)Bu(s)\,\mathrm{d}s\in X_{-1}$, where $\mathbb{T}$ is the semigroup generated by the operator $A$ from Assumption 2.1, and the output in the sense of distributions as $y(t)=\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\widetilde{W}_{C}\left[\begin{smallmatrix}\left(\mathcal{H}z\right)(b,t)\\ \left(\mathcal{H}z\right)(a,s)\end{smallmatrix}\right]$ where $z(t):=\int_{0}^{t}(t-s)x(t)\,\mathrm{d}s$.
@@
    By [21, Thm. 3.5], any generalised solution $(u,x,y)$ of a BIBO stable pHS with $u\in L^{\infty}_{\mathrm{loc}}\left([0,\infty),U\right)$ satisfies $y\in L^{\infty}_{\mathrm{loc}}\left([0,\infty),Y\right)$ (instead of only being a distribution) and for any $t>0$ we have that $\|y\|_{L^{\infty}\left([0,t],Y\right)}\leq c\|u\|_{L^{\infty}\left([0,t],U\right)}$, i.e. the inequality from Definition 2.3 extends to generalised solutions with bounded inputs.
The proposition below shows that the question of BIBO stability of pHS is not trivial.
Proposition 2.5.
@@
    There are BIBO stable pHS satisfying Assumption 2.1.
  2. (2)
    There are pHS satisfying Assumption 2.1, Equation (2.2) and $P_{0}^{*}=-P_{0}$ that are not BIBO stable.
@@
    Take $\textrm{pH}_{1}\left(1,0,1,\left[\begin{smallmatrix}1&0\end{smallmatrix}\right],\left[\begin{smallmatrix}0&1\end{smallmatrix}\right]\right)$. Then the system is just the trivially BIBO stable transport equation with input $u$ and output $y$
@@
    Take $\textrm{pH}_{1}\left(1,0,1,\left[\begin{smallmatrix}1&-1\end{smallmatrix}\right],\left[\begin{smallmatrix}0&1\end{smallmatrix}\right]\right)$. Then $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}=0$ and thus the system satisfies Assumption 2.1 by Remark 2.2 (and the semigroup is even contractive). On the other hand, the constant input $u(t)=1\in L^{\infty}\left([0,\infty),\mathbb{R}\right)$, yields the output $y(t)=\lfloor t\rfloor$ and hence the system is not BIBO stable.

### v2

# BIBO stability of 1-D hyperbolic boundary control systems

@@ Abstract.
We study the question of bounded-input bounded-output (BIBO) stability of a class of 1-D hyperbolic boundary control systems, which, in particular, contains distributed port-Hamiltonian systems. Exploiting the particular structure of the transfer function of these systems, we derive several sufficient conditions for BIBO stability.
Key words and phrases:
BIBO stability, port-Hamiltonian system, infinite-dimensional system, input-output stability
1991 Mathematics Subject Classification:
93D25, 93C05, 93C20, 35L04
@@ 1. Introduction
Spatially one-dimensional, hyperbolic evolution equations, such as given by port-Hamiltonian systems (pHS) [39, 21, 20] or hyperbolic balance laws [5, 11, 31, 3] have constituted an important system class used in modelling and control of real world applications. Their close connection to the ideas of energy-flow and energy-conservation makes them a useful framework for the study of physical systems[37, 38, 12] both mathematically as well as from an engineering perspective.
A particular type of example are systems of port-Hamiltonian type, as originally defined in [21, 40, 22, 19], and including those being impedance passive [2, 20, 37], which nowadays are commonly referred to as port-Hamiltonian systems. These include systems given by a PDE of the form
@@
on some spatial interval $[a,b]$ together with boundary control and observation at its endpoints. The justification and motivation for the study of theses systems lies in their ability to provide a modelling framework for a variety of different real-world processes, including first of all transport and flow phenomena such as in fluid mechanics [26] or chemistry [45, 44] and also vibration effects within flexible structures as appearing e.g. in robotics [24, 23]. Particularly, in this latter case, the approach using infinite-dimensional systems allows to refine finite-dimensional models where the mentioned phenomena (vibrations, flows, etc.) are usually ignored or only treated in a very simplified way. We refer the interested reader to the literature review [27] for a more history, and applications of distributed pHS. On the other hand, 1D hyperbolic systems have been extensively studied in the past decades, driven by numerous occurence of transport phenomena in the sciences, [6, 5, 16], such as in (traffic) networks, flows in fluid channels and light propagation in optics. In particular, stabilizing such systems is of key relevance both theoretically and practically.
@@
A natural and well-studied question for hyperbolic balance laws is whether a system is stable or, as often in a control context, whether it is stabilizable, see e.g. [6, 5] and the references therein. Only recently, work has been undertaken to systematically classify different notions for the above-mentioned systems of port-Hamiltonian type. These studies include internal stability, and in particular recent characterizations of exponential [34], asymptotic stability [42] and semi-uniform stability [1]. Previously only sufficient conditions for these properties were known [21, 40], which were however mainly restricted to the case of systems featuring a contraction semigroup as encountered in the vast majority of concrete examples.
@@
On the other hand, the input-output behaviour of boundary control systems is an important aspect for control applications. The systems considered here fall – under practically weak technical assumptions – in the class of $L^{2}$-well-posed systems, see e.g. [21]. Thus there exist constants $c_{t}$ for any $t>0$ such that for any input $u\in L^{2}\left([0,t],U\right)$ the corresponding output $y$ satisfies $\|y\|_{L^{2}\left([0,t],Y\right)}\leq c_{t}\|u\|_{L^{2}\left([0,t],U\right)}$. The importance of result like this is grounded in the port-Hamiltonian structure being used to model energy-flows and the input-output stability thus being related to the concepts of energy-conserving and energy-dissipating systems.
@@
In this contribution we study the notion of bounded-input bounded-output stability (BIBO stability), which is classical for finite-dimensional linear systems, for a class of 1-D hyperbolic boundary control systems. An input-output system is called BIBO stable if any bounded input is mapped to a bounded output and there exists a uniform relative norm bound [36, 9, 41, 10, 4, 30, 15]. We thus aim at input-output relations of the form $\|y\|_{L^{\infty}\left([0,t],Y\right)}\leq c\|u\|_{L^{\infty}\left([0,t],U\right)}$ for some positive constant $c$.
@@
The question of whether a system is BIBO stable or not, while relatively straightforward in the finite-dimensional setting, becomes significantly more involved in the case of an infinite-dimensional state space and control/observation acting on the spatial domain’s boundary. For instance, it is then no longer the case that BIBO stability is always implied by exponential stability (see [30, Thm. 5.1], but note that crucially this example is not hyperbolic). In [30] it was shown that, for a linear system described in the framework of system nodes (see [33]) with finite-dimensional input and output spaces – as will be the case in the system class considered in this contribution – BIBO stability is equivalent to the inverse Laplace transform of the transfer function $\mathbf{G}(s)$ being a measure of bounded total variation. This includes in particular the case that the impulse response – upon being well-defined – of the system is an $L^{1}$ function. While this condition offers a complete characterisation of when a system is BIBO stable, it is often of limited applicability for concrete systems. Typically no closed form expression for the inverse Laplace transform of the respective transfer function exists or is prohibitively hard to find. Thus BIBO stability is in general much more difficult to establish than e.g. $L^{2}$-well-posedness. This circumstance motivates the search for other sufficient or necessary conditions and the concrete study of BIBO stability for particular system classes. A few such results have been provided in for example [30] already, however almost exclusively for parabolic systems, which exclude the hyperbolic systems considered here.
@@
We note that BIBO stability – besides being an intriguing topic in itself and classical control theory – plays an essential role in e.g. the applicability of funnel control to relative degree systems [17, 7] and more generally for systems for which high-gain control is used. Such systems with interior dynamics modelled by port-Hamiltonian systems appear for example in the study of mechanical systems containing flexible and deformable parts [24], which significantly influence the behaviour.
@@
This paper begins with the definition of the considered class of hyperbolic 1-D boundary control systems, which is based on the original framework used for port-Hamiltonian systems in [21, 46] in Section 2. Then, in Section 3, we derive an abstract decomposition of the transfer function , which takes inspiration from recent results for exponential stability of pHS [34] and exploit its particular structure in order to apply the known characterisation of BIBO stability in a general setting. This yields a collection of sufficient criteria expressed in the form of matrix conditions for a subclass of pHS. Section 4 then presents a different approach to BIBO without any reference to transfer functions, instead employing the $L^{1}$-well-posedness of pHS on an $L^{1}$ state space and the connection between $L^{1}$ and $L^{\infty}$ stability notions discussed in [29].
@@ 1.1. Notation
Let $A$ be a linear operator defined on some Banach space $X$. Then we denote by $\mathrm{D}(A)\subseteq X$ its domain of definition, by $\rho(A)$ its resolvent set and by $\sigma(A)$ its spectrum. If $A$ is further the generator of a strongly continuous semigroup $\mathbb{T}=\left(\mathbb{T}(t)\right)_{t\geq 0}$ on $X$, then $X_{1}$ is defined as the space $\mathrm{D}(A)$ with the norm $\|x\|_{X_{1}}:=\|(\beta I-A)x\|_{X}$ with $\beta\in\rho(A)$ and $X_{-1}$ as the completion of $X$ with respect to the norm $\|x\|_{X_{-1}}:=\|(\beta I-A)^{-1}x\|_{X}$ again with $\beta\in\rho(A)$. There exists a unique extension $\mathbb{T}_{-1}=\left(\mathbb{T}_{-1}(t)\right)_{t\geq 0}$ of the semigroup $\mathbb{T}$ to the space $X_{-1}$ with generator $A_{-1}:X\rightarrow X_{-1}$ which is an extension of the operator $A$. For more details see [33, 35].
Let $\mathbbm{1}_{n}\in\mathbb{K}^{n\times n}$ denote the identity matrix. For a matrix $M\in\mathbb{K}^{n\times n}$ and $p,q\in[1,\infty]$ let $\|M\|_{\ell^{p}\rightarrow\ell^{q}}$ be the matrix norm induced by the $\ell^{p}$-norms on $\mathbb{K}^{n}$ given by $\|x\|_{\ell^{p}}=\left(\sum_{k=1}^{n}|x_{k}|^{p}\right)^{\frac{1}{p}}$ for $p\in[1,\infty)$ and $\|x\|_{\ell^{\infty}}=\sup_{1\leq k\leq n}|x_{k}|$ for $p=\infty$ as
@@
For a Borel measure $h$ on $0,\infty)$, let $\mathcal{L}\left\{h\right\}$ denote its Laplace transform if it exists (on some right half-plane). For a function $g$ of one complex variable, if there exists a measure $\mu$ such that $\mathcal{L}\left\{\mu\right\}$ and $g$ agree on some right half plane we will call $\mu$ the inverse Laplace transform and write $\mu=\mathcal{L}^{-1}\left\{g\right\}$.
Let $\mathcal{M}([0,\infty),\mathbb{K}^{n\times m})$ denote the set of Borel measures of bounded total variation on $0,\infty)$ with values in $\mathbb{K}^{n\times m}$. Furthermore, for $h\in\mathcal{M}([0,\infty),\mathbb{K}^{n\times m})$ let $\|h\|_{\mathcal{M}}$ denote the total variation of $h$ [14, Sec. 3.2].
Furthermore, let $M_{|\cdot|}\in\mathbb{K}^{n\times n}$ be defined as the matrix with entries $(M_{|\cdot|})_{ij}=|M_{ij}|$.
2. A class of hyperbolic boundary control systems
In this work we consider linear hyperbolic boundary control systems of the form
(1) $\frac{\partial x}{\partial t}(\xi,t)$ $=P_{1}\frac{\partial}{\partial\xi}\left(\mathcal{H}(\xi)x(\xi,t)\right)+P_{0}(\xi)\left(\mathcal{H}(\xi)x(\xi,t)\right),$ $(\xi,t)$ $\in[a,b]\times[0,\infty),$
@@ 2. A class of hyperbolic boundary control systems
with the following assumptions:
Assumption 1.
  1. (1)
@@
  3. (3)
    $\mathcal{H}\in L^{\infty}\left([a,b],\mathbb{K}^{n\times n}\right)$, $\mathcal{H}(\xi)$ is self-adjoint for almost all $\xi\in[a,b]$ and $\exists M,m>0$ s.t. $m\mathbbm{1}_{n}\leq\mathcal{H}(\xi)\leq M\mathbbm{1}_{n}$ for almost all $\xi\in[a,b]$;
  4. (4)
    $\widetilde{W}_{B}\in\mathbb{K}^{n\times 2n}$ and $\left[\begin{smallmatrix}\widetilde{W}_{B}\\ \widetilde{W}_{C}\end{smallmatrix}\right]\in\mathbb{K}^{2n\times 2n}$ have full rank;
  5. (5)
@@
    is the infinitesimal generator of a $C_{0}$-semigroup on the state space $X=L^{2}\left([a,b],\mathbb{K}^{n}\right)$ with inner product
    $\left\langle f,g\right\rangle_{X}:=\frac{1}{2}\int_{a}^{b}g(\xi)^{*}\mathcal{H}(\xi)f(\xi)\,\mathrm{d}\xi.$
We will in the following use $\Sigma_{n}(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C})$ to denote this hyperbolic boundary control system (hBCS).
Note that a similar class was studied in the seminal work [22, 46] in the context of $L^{2}$-well-posedness, which marked the beginning of what later became known as distributed port-Hamiltonian systems [21]. We emphasise that the class studied here particularly includes these port-Hamiltonian systems, both with [2, 20] or without [21] assumed impedance passivity, which in the meantime seems to be commonly assumed for port-Hamiltonian systems. Our notation used here originates from the text book [21] and as there we will refer to $\mathcal{H}$ as the Hamiltonian of the system.
Remark 2.
@@
    Since $A$ is a bounded perturbation of the operator $\frac{\partial}{\partial\xi}\mathcal{H}$, with the same domain, the assumption that $A$ generates a $C_{0}$-semigroup, Assumption 1(5), is independent of the choice of $P_{0}$.
@@
    The Assumption 1(5) of AA generating a $C_{0}$-semigroup is in particular satisfied if
    (2) $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}\geq 0$
    where $R_{0}=\frac{1}{\sqrt{2}}\left[\begin{smallmatrix}P_{1}&-P_{1}\\ \mathbbm{1}_{n}&\mathbbm{1}_{n}\end{smallmatrix}\right]$ and $\Sigma=\left[\begin{smallmatrix}0&\mathbbm{1}_{n}\\ \mathbbm{1}_{n}&0\end{smallmatrix}\right]$. If in addition $P_{0}^{\ast}=-P_{0}$, then (2) is even equivalent to the semigroup being contractive [21, Thm. 7.2.4, Thm. 10.3.1].
For our system class, direct extension of [21, Thm. 11.3.2 & Thm. 12.1.3] to the case of spatially dependent $P_{0}$, shows that the transfer function $\mathbf{G}$ can be found for $s\in\rho(A)$ as the unique solution of
[…]
2.1. BIBO stability of hyperbolic boundary control systems
In this paper we want to study the question of BIBO stability of systems (1) satisfying Assumption 1. In the following we will understand the term classical solution in the common sense of evolution equations, that is as a collection $(u,x,y)$ with $u\in C([0,\infty),U)$, $x\in C^{1}([0,\infty),X)$ and $y\in C([0,\infty),Y)$ solving Equation (1).
Definition 3.
A hBCS is called BIBO stable if there exists $c>0$ such that for any of its classical solution $(u,x,y)$ with $u\in C^{\infty}_{c}([0,\infty),U)$ and $x(0)=0$ and for all $t>0$ we have that $\|y\|_{L^{\infty}\left([0,t],Y\right)}\leq c\|u\|_{L^{\infty}\left([0,t],U\right)}$.
Remark 4.
Any hBCS satisfying Assumption 1 can be equivalently expressed within the framework of system nodes (see [33] for a general introduction and [21, Ch. 11][28] for the rewriting of hBCS respectively general boundary control systems in this form). In particular, the following observations can be made.
@@ 2.1. BIBO stability of hyperbolic boundary control systems
    Following the construction in [33], we can define a generalised solution $(u,x,y)$ of a hBCS for any input function $u\in L^{1}_{\mathrm{loc}}\left([0,\infty),U\right)$ and initial value $x_{0}\in X$.
    Here the state is given by $x(t)=\mathbb{T}(t)x_{0}+\int_{0}^{t}\mathbb{T}_{-1}\left(t-s\right)Bu(s)\,\mathrm{d}s\in X_{-1}$, where $\mathbb{T}$ is the semigroup generated by the operator $A$ from Assumption 1(5), and the output in the sense of distributions as $y(t)=\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\widetilde{W}_{C}\left[\begin{smallmatrix}\left(\mathcal{H}z\right)(b,t)\\ \left(\mathcal{H}z\right)(a,s)\end{smallmatrix}\right]$ where $z(t):=\int_{0}^{t}(t-s)x(t)\,\mathrm{d}s$.
@@
    The hBCS considered here fall within the class of systems investigated for BIBO stability in [30]. In particular, by [30, Thm. 3.5], any generalised solution $(u,x,y)$ of a BIBO stable hBCS with $u\in L^{\infty}_{\mathrm{loc}}\left([0,\infty),U\right)$ satisfies $y\in L^{\infty}_{\mathrm{loc}}\left([0,\infty),Y\right)$ (instead of only being a distribution) and for any $t>0$ we have that $\|y\|_{L^{\infty}\left([0,t],Y\right)}\leq c\|u\|_{L^{\infty}\left([0,t],U\right)}$, i.e. the inequality from Definition 3 extends to generalised solutions with bounded inputs.
That the question of BIBO stability of the system class considered here is not trivial in either way is shown by the following proposition.
Proposition 5.
@@
    There exists a BIBO stable hBCS satisfying Assumption 1.
@@
    There exists a hBCS $\Sigma_{n}(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C})$ satisfying Assumption 1, $P_{0}^{*}=-P_{0}$ and such that Equation (2) holds, which is not BIBO stable.
@@
    Take $\Sigma_{1}\left(1,0,1,\left[\begin{smallmatrix}1&0\end{smallmatrix}\right],\left[\begin{smallmatrix}0&1\end{smallmatrix}\right]\right)$. Then the system is just the trivially BIBO stable transport equation with input $u$ and output $y$
@@
    Take $\Sigma_{1}\left(1,0,1,\left[\begin{smallmatrix}1&-1\end{smallmatrix}\right],\left[\begin{smallmatrix}0&1\end{smallmatrix}\right]\right)$. Then $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}=0$ and thus the system satisfies in particular Assumption 1(5) by Remark 2 (and the semigroup is even contractive). On the other hand, the constant input $u(t)=1\in L^{\infty}\left([0,\infty),\mathbb{R}\right)$, yields the output $y(t)=\lfloor t\rfloor$ and hence the system is not BIBO stable. ∎
It turns out that even restricting to a subtype of the system class, i.e. adding further assumptions on the boundary control systems, such as impedance passivity (see e.g. [2, Sec. 3.2]), does not improve the situation.
Example 6 (Impedance passive but non-BIBO stable pHS).
Consider the system $\Sigma_{2}\left(\mathbbm{1}_{2},0,\left[\begin{smallmatrix}\frac{1}{2}&0\\ 0&1\end{smallmatrix}\right],\widetilde{W}_{B},\widetilde{W}_{C}\right)$ on $[0,1]$ with
$\widetilde{W}_{B}=\begin{bmatrix}-1&0&\frac{1}{2}&\frac{1}{2}\\ 0&-1&-\frac{1}{2}&-\frac{1}{2}\end{bmatrix}\quad\textrm{and}\quad\widetilde{W}_{C}=\frac{1}{2}\begin{bmatrix}-1&-1&-1&-1\\ 1&-3&1&-3\end{bmatrix}$,
that is the system given by the equations
$\frac{\partial x_{1}}{\partial t}(\xi,t)$ $=\frac{1}{2}\frac{\partial x_{1}}{\partial\xi}(\xi,t),$ $(\xi,t)$ $\in[a,b]\times[0,\infty),$
[…]
$u_{1}(t)$ $=-\frac{1}{2}x_{1}(1,t)+\frac{1}{4}x_{1}(0,t)+\frac{1}{2}x_{2}(0,t),$ $t$ $\in[0,\infty),$
[…]
$y_{1}(t)$ $=-\frac{1}{4}x_{1}(1,t)-\frac{1}{2}x_{2}(1,t)-\frac{1}{4}x_{1}(0,t)-\frac{1}{2}x_{2}(0,t),$ $t$ $\in[0,\infty),$
$y_{2}(t)$ $=\hphantom{-}\frac{1}{4}x_{1}(1,t)-\frac{3}{2}x_{2}(1,t)+\frac{1}{4}x_{1}(0,t)-\frac{3}{2}x_{2}(0,t),$ $t$ $\in[0,\infty).$
Using the characterisation from [2, Prop. 3.2.16] we can show that this is an impedance passive port-Hamiltonian system.
However, this system is up to the choice of $\widetilde{W}_{C}$ the same as the one shown to be not BIBO stable in Proposition 7. As $\left[\begin{smallmatrix}\widetilde{W}_{B}\\ \widetilde{W}_{C}\end{smallmatrix}\right]$ has full rank, we can write the outputs $y_{1}$ and $y_{2}$ as linear combinations of the inputs and outputs of that system. Hence this system cannot be BIBO stable either.